OpenStudy (anonymous):

4. Express the series below in summation notation for the specified number of terms. b. 10 + 14 + 18 + 22

OpenStudy (solomonzelman):

each next term is +4 series starts from "10" there are 4 terms in the series

OpenStudy (anonymous):

so how would we put that into the equation

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \sum_{n=10}^{4} ~~4(\color{red} { n } -1)+10 } \]

OpenStudy (solomonzelman):

red n is nth term, like 2nd term 2rd term....

OpenStudy (anonymous):

whata bout 1/2+2/3+3/4+4/5+5/6

OpenStudy (solomonzelman):

so for 1st term you would put in 1 1st term = 4(1-1) + 10 = 0 + 10 = 10 2nd term = 4(2-1) + 10 = 4 + 10 = 14 3rd term = 4(3-1) + 10 = 8 + 10 = 18 4th term = 4(4-1) + 10 = 12 + 10 = 22 I don't get the last you thing you wrote

OpenStudy (anonymous):

4. Express the series below in summation notation for the specified number of terms. \[\frac{ 1 }{ 2 }+\frac{ 2 }{ 3}+\frac{ 3 }{ 4 }+\frac{ 4 }{ 5 }+\frac{ 5 }{ 6 }\]

OpenStudy (solomonzelman):

\[\huge\color{green}{ \sum_{n=1}^{5} ~~\frac{n}{n+1} }\]

OpenStudy (solomonzelman):

How does this one look :) ?

OpenStudy (anonymous):

good

OpenStudy (anonymous):

whata bout this

OpenStudy (anonymous):

Express the series below in summation notation for the specified number of terms. 9. 0 + 3 + 8 + 15 + 24 + 35 + 48 + 63 + 80

OpenStudy (solomonzelman):

+3, +5, +7, +9, +11 .... \[\huge\color{blue}{ \sum_{n=2}^{9} ~~(2n+1) }\]